OpenStudy (anonymous):

If y = 5x – 4, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? (4 points) {(0, 4), (2, –6), (4, 20)} {(0, –4), (2, 6), (4, 16)} {(0, –4), (2, 6), (4, 20)} {(0, 4), (2, 6), (4, 16)}

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (anonymous):

A point is (x,y). So when you plug in each x and simplify, you should get the y-value from the same point. If you do, then that's a good input/output pair. If not, then move on to the next answer. For example, to test the first point: (0, 4) --> 5(0) + 4 = 4, which is what we should get, so this point works. Now you have to try the other two points in that row.

OpenStudy (anonymous):

okay... I still dont get it lol sorry

OpenStudy (jordanloveangel):

Each coordinate is (x,y). You want plug in the x values {0,2,4} to figure out the corresponding y values.

OpenStudy (jordanloveangel):

For example, for x=0: (x,y)=(x,5x−4)=(0,5(0)−4)=(0,−4) So one point is (0,−4).

OpenStudy (jordanloveangel):

This eliminates all but: {(0,−4),(2,6),(4,16)}{(0,−4),(2,6),(4,20)}

OpenStudy (jordanloveangel):

Okay, first step: We know that y=5x−4, right? This means wherever there is a y, we can put 5x−4 and it will still be equal. So: (x,y)=(x,5x−4) Does this make sense?

OpenStudy (anonymous):

yes...

OpenStudy (jordanloveangel):

okay glad medal please

OpenStudy (anonymous):

so would it be C?

OpenStudy (jordanloveangel):

yes...

OpenStudy (anonymous):

thanks!

OpenStudy (jordanloveangel):

yeah np